OpenStudy (anonymous):

solve the equation 30 = -3 (x + 4) + 2x

OpenStudy (anonymous):

-42 = x

OpenStudy (anonymous):

I think...

OpenStudy (anonymous):

x=-25

OpenStudy (anonymous):

if you wrote it write than it would be impossible because there would a a negative x are you were this is the right problem

OpenStudy (anonymous):

haley_ann1396 is correct

OpenStudy (anonymous):

x=-42

OpenStudy (anonymous):

30 = -3(x+4) +2x 30 = -3x - 12 + 2x 30 = -x - 12 42 = -x -42 = x

OpenStudy (anonymous):

yes haley_ann1396 is right